OpenStudy (anonymous):

Find the slope of line passing through the points (5, 9) and (5, -2)

OpenStudy (danjs):

hello, welcome to open study!

OpenStudy (danjs):

do you recall the slope formula slope = difference in y / difference in x

OpenStudy (danjs):

\[slope = \frac{ y - y1 }{ x-x1 }\] where the two points are (x,y) = (5, 9) (x1,y1) = (5, -2) plug in the values, and calculate. .

OpenStudy (danjs):

@emilypfranta

OpenStudy (anonymous):

solve it @emilypfranta

OpenStudy (anonymous):

i got 11

OpenStudy (danjs):

(-2-9) / (5-5) = -11 / 0 Do you know what dividing by zero means

OpenStudy (danjs):

The slope is undefined here, or you could say infinity i guess. It is a VERTICAL line. Notice the X values are the same for both points, 5

OpenStudy (danjs):

Here is Your line connecting those 2 points

OpenStudy (anonymous):

Okay i had my answer as undefined at first, but i wasn't sure how to solve it. Now i understand.

OpenStudy (danjs):

that is one of the special cases.. #/0 = Slope Undefined, Vertical Line 0/# = Zero Slope, Horizontal line